Mathematics 119, Fall Semester 2000-2001

Are you in the right course?

Math 119-120 is intended for students planning to enter pre-professional or biology programs. It is not intended for students who intend to major in engineering, physics, mathematics, or most of the chemistry programs. Those majors require Math 125-126.

Instructor:  Professor Joachim Rosenthal

309 CCMB

Phone: 631-8370, e-mail: Rosenthal.1@nd.edu

Office Hours:  Tuesday 9:00–10:00

Friday 9:00–10:00 Or by appointment.

Tutorial Instructor: Maria Kiskowski

222 CCMB, e-mail: Kiskowski.1@nd.edu

TA’s Office Hours: Tuesday 2:00–3:00 or by appointment.
Text:  Single Variable Calculus (fourth edition),

by James Stewart

 Course Web Page: http://www.nd.edu/∼rosen/math119

Classes, Tutorials and Help Sessions

Class meets at 11:45 MWF in Pasquerilla Center 112. Students are required to attend these classes.

Each student is also assigned to a Thursday tutorial section. It is your responsibility to know the time, place and section number of your tutorial section. At the tutorial, the previous week’s homework will be returned. Students will then be encouraged to ask questions about that homework or about the current homework. There will be sometimes quizzes given during tutorial time. The purpose of the tutorial is to help students master the material currently being covered.

In addition to the tutorials, the professor and the tutorial instructor will have office hours each week at which you can get assistance in understanding the course work and doing the homework problems.

Examinations, homework and grades

There will be three one-hour examinations and one final examination (whose dates, times and locations are listed below). Each one-hour exam will be worth 100 points, and will be returned to you at the following tutorial session. The final exam is a two-hour exam and will be worth 150 points. The final exam will cover all the material of the course. There will be a total of 25 points for the quizzes and a total of 25 points for the homework. There are a total of 500 possible points for the semester, and the cut-offs for the semester grades will be based on this total.

Homework will be due at the Thursday tutorial and returned the following week. Usually three assignments will be due each Thursday; specific assignments due each week will be announced in that week’s lectures. Because it may happen that you have trouble with some homework problems and want another shot at them after you see the TA, we will accept homework as late as 4:00 of the following day (Friday). If you do not turn it in to the TA at the tutorial, you have to come to the Mathematics Department and put your homework in her mailbox (on the 3rd floor of CCMB).

The main purpose of collecting and returning homework is to let you know if you are doing the problems correctly. The homework grade is designed to reward effort. Each problem is graded either 0 (if missing or complete nonsense) or 1 (for any honest attempt). So the total number of points on any assignment is simply the number of problems honestly attempted.

All examinations and homework are conducted under the honor code. While cooperation in doing homework is permitted (and encouraged), copying is not. Exams are closed book and are to be done completely by yourself with no help from others. Basic graphing calculators (like TI 82/83 and TI 85/86) are allowed on the exams. (No Palmtops or Labtops however!)

A student who misses an examination will receive zero points for that exam unless he or she has written permission from the Dean of the First Year of Studies. (An excuse is almost certainly not going to be accepted if it is presented after the exam takes place.) Please be aware that travel plans are not considered to be a valid excuse by the Dean of the First year of Studies.

Exams

Exam 1 Thursday, September 21, 2000 8:00 AM DBRT 138
Exam 2 Thursday, October 26, 2000  8:00 AM DBRT 138
Exam 3 Tuesday, November 28, 2000 8:00 AM DBRT 138
Final Friday, December 15, 2000 1:45 PM

(Location of final exam will be announced later.)

Mathematics 119, Fall Semester 2000-2001 HOMEWORK ASSIGNMENTS

Number Assignment
1  p. A15/A16 # 9, 10, 23–28, 31–33, 37–40
2  p. 22/23 # 1–2, 5–6, 18, 23–26, 33–36
3  p. A15  # 2–5
p. A23  # 1–8
4  p. 71/72 # 2–6
5  p. 82/83 # 7–9, 12–14
6 p. 92 # 3–4, 11–14, 19, 21, 23
7  p. 112/113 # 3–4, 13–14, 29–30, 34–35
8  p. 122  # 3, 7–9, 17
p. 134  # 3–5, 7–8
9  p. 134  # 13–15, 23, 25–26
p. 145/146 # 15–18, 30
10  p. 156/157 # 1, 2, 4, 10–13, 23, 24, 34, 36–39, 59, 60
11  p. A32/A33 # 1–2, 7–8, 13–15, 23, 27, 28, 30, 32, 42–46, 53, 65, 66, 70
12  p. 167/168 # 12–13, 16
p. 175  # 1, 3, 5–8, 10–14, 16, 21
13  p. 176  # 25–26, 29–30, 32, 35–38
14  p. 183  # 7, 9–11, 13–14, 17, 33–36, 38, 49, 51–53
15  p. 190  # 6–7, 9, 13, 14, 16, 25
16  p. 197  # 5–6, 16–18, 25–27, 29, 41
17  p. 168  # 17–19
p. 198  # 45–46, 49, 51
18  p. 203  # 1, 3, 5–10
19  p. 204  # 11, 15–20, 25
20  p. 231  # 19–24, 31–33, 35–36, 47–51, 67
21  p. 238/239 # 1–4, 11–14, 17
22  p. 247/248 # 5–8, 13–15, 17–18, 27–30
23  p. 261  # 7–8, 11–12, 23–24, 39
24  p. 270  # 4–5, 15–16, 39–40
25  p. 276  # 1, 3, 6–8
26  p. 283  # 9–15, 19
27  p. 304  # 2, 3–8, 29–32
28  p. 322/323 # 3–4, 11, 15–17
29  p. 334/335 # 2–3, 8, 29
30  p. 344/345 # 5–8, 17–20, 49–50
31  p. 352/353 # 2–3, 7–8, 17–20, 46
32  p. 361/362 # 1–4, 7–10, 60
33  p. 376  # 5, 16–18, 25

Learning Resource Center(LRC) Programs for Mathematics Courses Fall 2000

The LRC is a part of the First Year of Studies. Among the many services provided to first year students, it offers the Tutoring Program, the Collaborative Learning Program, and a series of Math Workshops. The purpose of these programs is to assist the students with their proficiency in Mathematics.

Tutoring Program: Tutoring is available to students whose success in a first year course requires assistance beyond that which can be given by the teacher or staff of the courses. Tutoring sessions for small groups of students are offered in Math 104, 105, 108, 111, 119, 120, 125, 126, 225, 228. These sessions meet once weekly for two hours and will begin Tuesday, September 5, 2000.

Collaborative Learning Program: Small groups of students (4-6) are grouped by course and section, and work on mathematics homework. In the process of talking with each other and sharing their ideas about the work, they learn more Mathematics and understand it better. Sometimes groups share their ideas with other groups that are working on the same problems. In this way, all the students gain valuable ideas from others. Facilitators, upper class students who have successfully completed the course, are available to encourage problem solving and collaboration among group members, and to help when no one in the group can answer a question. Math 105, 119, 125 groups meet twice weekly for one and one-half hours each time and Math 104, 108, 120, 126 groups meet once weekly for two hours, all in the main room of the LRC, beginning Monday, September 11, 2000 with the following schedule:

Math 104 Monday  6:00-8:00 pm
Math 105 Sunday & Wednesday 6:00-7:30 pm
Math 108 Thursday  4:00-6:00 pm
Math 119 Tuesday & Thursday 8:30-10:00 pm
Math 120 to be announced
Math 125 Monday & Wednesday 4:00-5:30 pm OR Tuesday & Thursday 6:00-7:30 pm
Math 126 Sunday  8:30-10:30 pm

Workshops: Workshops are help sessions which do \underline{not} require regular attendance. Each session is held every week, beginning September 5, 2000, with the following schedule:

Math 104 Sunday 4:00-6:00 pm OSHA 203
Math 105 Sunday 6:00-8:00 pm OSHA 203
Math 108 Sunday 6:00-8:00 pm OSHA 203A
Math 119 Sunday 6:00-8:00 pm OSHA 206
Math 125 Monday 7:00-9:00 pm OSHA 207
Math 126 Sunday 6:00-8:00 pm OSHA 208

Sign-up Instructions: To sign up, please see Nahid Erfan, Director of the LRC. Notice, there is no need to sign up for Workshops.

Attendance Policy: Students are expected to attend all tutoring or collaborative learning sessions (but not all workshops). If unable to attend a session, call the LRC at 631-5294 in order to obtain an excused absence. Two absences may result in dismissal and readmission cannot be guaranteed.